# LASeR: Learning to Adaptively Select Reward Models with Multi-Armed Bandits

### v1

@@ 3.2 Bandit Algorithms for Adaptive RM Selection
The multi-armed bandit (MAB) problem addresses the challenge of balancing exploration and exploitation in sequential decision-making (Vermorel & Mohri, 2005; Audibert et al., 2009). The goal is to maximize cumulative MAB rewards over time by selecting arms that yield the highest MAB rewards. [^3] In order to distinguish between rewards or scores generated by RMs and the rewards used in MAB literature, we refer to the latter as “MAB rewards”. A decision-making agent faces a trade-off: whether to exploit the arm with the highest known MAB reward based on past observations or explore other, less familiar arms to gather more information that might lead to even better rewards in the future. In a contextual MAB setting, the agent is also provided with additional information in the form of a context, such as current state and input, to help inform arm selection accordingly.
RM selection in LASeR.
@@
LASeR uses mini-batch training for each iteration, i.e., we use MABs to select a single RM for a batch of prompts $\mathbf{x}_{t}$ for $t^{\text{th}}$ batch or training step (total of $T$ steps/batches in each iteration). [^4] Note that LASeR can switch between RMs at the instance level if the batch size is set to 1; however, for the sake of efficiency, we batch instances together both for LASeR and the baselines, as this reduces the computational overhead associated with loading RMs onto the GPU. Let the set of $K$ reward models (or arms) denoted by =$\mathcal{R}=\{R_{1},R_{2},\dots,R_{K}\}$, where each $R_{k}$ corresponds to a different RM. We employ LinUCB (Li et al., 2010), a contextual bandit algorithm for the arm or RM selection. Li et al. (2010) assume that the MAB reward – in our case, the cumulative train loss function on the batch ($\mathcal{L}^{m}$) at the given iteration $m$ – can be modeled linearly as a function of context features and computes the expected MAB reward of each arm with an upper confidence bound to ensure exploration (Garivier & Moulines, 2008; 2011). In each step $t$, we have a batch of input prompts $\mathbf{x}_{t}$ for which we compute sentence embeddings, using the policy model $\pi_{m}$, and use the mean sentence embedding as the context $c(t)$ to the MAB, i.e., $c(t)=\sum_{x\in\mathbf{x}_{t}}e_{m}(x)/|\mathbf{x}_{t}|$, where $e_{m}(.)\in\mathbb{R}^{d}$ yields the sentence embedding from the model $\pi_{m}$. We calculate the embedding for a prompt as the last token embedding from model $\pi_{m}$. The learned parameters of the LinUCB bandit include $\hat{\theta}_{k}\in\mathbb{R}^{d}$ which represents the learned weights for the features of each reward model and $A_{k}\in\mathbb{R}^{d\times d}$ (a covariance matrix) and a bias vector $b_{k}\in\mathbb{R}^{d}$ corresponding to each arm or RM $R_{k}$. We initialize the parameters for LinUCB by randomly initializing $b_{k}$ and setting parameter $A_{k}$ to the identity matrix. Based on the LinUCB algorithm, for each batch, the selected RM $R^{\star}_{t}$ is determined by:
$R^{\star}_{t}=\arg\max_{k\in[1,K]}\left(c(t)^{\top}\hat{\theta}_{k}+\alpha\sqrt{c(t)^{\top}A_{k}^{-1}c(t)}\right),$ (2)
@@ 3.3 LLM and Bandit Training in LASeR
LASeR with Best-of-$n$ Sampling.
For settings where finetuning the LLM is not desirable or feasible, LASeR can also be applied to learn the MAB parameters without training the LLM. Rather than fine-tuning the model with preference data, we employ best-of-$n$ sampling (Lightman et al., 2023; Sun et al., 2024), where multiple responses are generated, and the best one is selected based on the RM. The bandit parameters are then updated using equation (3), with the MAB reward calculated as the negative normalized NLL loss on the train data. This updated bandit can subsequently be used for inference on the test set. This approach is particularly useful for tasks such as long-context understanding, where training would be too computationally intensive (example setting in Sec. 4.2).
@@ 4.1 Experimental Setup
We conduct our experiments on the Llama-3-8B base (AI@Meta, 2024) and the Mistral-7b-v3-instruct (Jiang et al., 2023) models. For training, all models are fine-tuned using Low-Rank Adaptation (LoRA) (Hu et al., 2021) for efficiency. For both training and inference, we do 0-shot prompting and sampling $n\!=\!30$ responses per prompt with temperature $0.8$ (see Sec. A.1).
Reward models.
@@
We select $K=4$ strong 7B RMs from RewardBench (Lambert et al., 2024), which include Zephyr-7B-Alpha, Qwen1.5-7B-Chat, Eurus-7B-KTO, and OLMo-7B-Instruct. Following the pipeline outlined in Lambert et al. (2024), for these models, we compute the reward for each response as the log likelihood of the RM for that response (details in Sec. A.1).
Datasets and Metrics.
Our experiments cover a range of tasks and datasets (see Sec. A.1):
@@
  • Reasoning: Evaluating reasoning abilities is crucial for testing the model’s capacity to handle complex, multi-step tasks and has presented a challenge to iterative preference optimization methods (Yuan et al., 2024; Chen et al., 2024b). We train and evaluate on StrategyQA (Geva et al., 2021), MMLU (Hendrycks et al., 2021b; a), and GSM8K (Cobbe et al., 2021).
@@
  • Instruction-Following: We further evaluate our method on heterogeneous tasks without gold labels. We use user prompts from WildChat dataset (Zhao et al., 2024), which contains a collection of natural user-chatbot interactions. This dataset has five primary categories of instruction-following prompts: creative writing, analysis, coding, factual information, and math reasoning. Due to computational constraints, we randomly subsample 5K prompts from each category for model training. We compare models trained with LASeR against baselines (described below) using length-controlled AlpacaEval (Dubois et al., 2024) that pairs responses from two different LLMs and uses GPT-4 as a judge to pick the winner, accounting for the length of both responses.
@@
  • Long-Context Understanding: As finetuning LLMs on long-context inputs is computationally intensive, we demonstrate the effectiveness of LASeR using Best-of-$n$ sampling on LongBench (Bai et al., 2023) which consists of multiple tasks, such as single-document QA, multi-document QA, summarization, and few-shot learning. For the QA and few-shot learning tasks, we measure performance with F1 score, while for summarization we use Rouge-L (Lin, 2004).
Baselines.
We compare our models against the following baselines:
@@
  • Best RM: From our collection of RMs, we pick the RM that corresponds to the best overall score on RewardBench (Lambert et al., 2024): Zephyr-7B-Alpha. We use this single RM during training (c.f. Sec. 3.1). This baseline reflects the performance gain a user could expect when selecting the best RM from a leaderboard without knowing a priori how it generalizes to a particular domain.
@@
  • Avg. RM: Here, we perform single RM training over all the RMs in our collection and report the average performance. A comparison with this baseline represents an expected gain from a randomly-picked RM from a leaderboard.
@@
  • Random RM Selection: In this baseline, we randomly sample a single RM from the set of RMs (from a uniform distribution) for each training batch in every iteration. This comparison demonstrates whether, without prior knowledge of which RM is best suited for a downstream task, LASeR can outperform random sampling of RMs during training.
@@
  • Sequential RM Selection: In training, this method explores different RM sequentially and based on a set order in each iteration to examine their impact on model training, demonstrating that, instead of optimizing with all RMs, LASeR can adaptively select the best RM for each batch.
@@
  • Offline RM Ensemble: We generate multiple responses for each query, which are scored (offline) using each RM, and the preference dataset is created by averaging the scores across all RMs (following Coste et al. (2023)); thus, comparing LASeR with using all RMs simultaneously.
Conceptually, the best RM baseline serves as an “exploit-only” setting that exploits the best available RM based on RewardBench. On the other hand, the random and sequential selection baselines are “explore-only” in that they pick a new RM either randomly or via a predefined sequence irrespective of the performance of each arm (RM). In our experiments, we train each model to convergence. In particular, LASeR, “Best RM”, “Avg. RM”, and RM ensemble baselines were trained for 10 iterations. For both the sequential and random RM selection, we found LLM training took longer to converge, and consequently, the model was trained for 25 iterations. The number of iterations for each approach was chosen based on stabilization in performance on the dev set (see Sec. A.1).
Table 1: Performance on reasoning benchmarks. The baselines also include supervised fine-tuning on human-written responses (SFT) as a reference for performance without preference optimization. The highest accuracy is shown in bold, and the second-highest accuracy is underlined. Across both Llama-3-8B and Mistral-7B models, LASeR yields the highest accuracy for each task.
Method Llama-3-8B Mistral-7B
StrategyQA GSM8K MMLU Avg. StrategyQA GSM8K MMLU Avg.
SFT 80.41 69.43 65.66 71.83 68.57 43.62 56.48 56.22
Best RM 84.29 73.16 67.15 74.87 70.06 45.81 62.04 59.30
Avg. RM 82.62 71.57 66.67 73.62 69.62 45.47 59.58 58.22
Offline RM Ensemble 82.96 70.94 67.04 73.65 68.89 44.53 58.23 57.22
Random RM Selection 84.37 71.99 67.85 74.74 69.97 45.12 59.88 58.32
Sequential RM Selection 83.90 72.94 68.02 74.95 70.59 46.11 59.66 58.79
LASeR (Ours) 85.96 74.75 68.24 76.32 73.06 46.89 62.27 60.74
@@ 4.2 Main Results
LASeR achieves the best average accuracy on reasoning tasks.
@@
Table 1 shows our method consistently outperforms the baselines across multiple benchmarks, particularly in the StrategyQA and GSM8K tasks. For example, using the Llama-3-8B model, LASeR yields the highest reasoning accuracy average across all tasks, with improvements of approximately $2\%$ absolute accuracy over the sequential baseline on both GSM8K and StrategyQA. On the Mistral-7B model, LASeR also improves average accuracy by roughly $2\%$ over the sequential baseline. Additionally, our method outperforms the best RM (based on RewardBench) baseline in average accuracy by $1.45\%$ and $1.44\%$ absolute accuracy with Llama-3-8B and Mistral-7B, respectively. In cases where the best RM is not known beforehand, LASeR surpasses the performance of the average RM by $2.7\%$ on Llama-3-8B and using RM ensemble for each instance by $2.67\%$ and $3.52\%$ on Llama-3-8B and Mistral-7B, respectively (in accuracy averaged over the three tasks). Overall, LASeR provides consistent results while the underlined second-place models show inconsistent performance across datasets and models. These results also emphasize the benefit of LASeR, as it eliminates the need to choose a different RM in advance, such as random, sequential selection, or the best RM.
LASeR beats baselines at instruction-following.
@@
Often, LLMs are used by large numbers of people with a diverse set of queries, goals, and intentions, and their preferences vary based on the underlying query. To demonstrate the effectiveness of LASeR in such settings, we compare the instruction-following performance, i.e., AlpacaEval win rates, of LLMs trained using LASeR with the baselines using WildChat (see Sec. 4.1). Specifically, with Llama-3-8B, LASeR achieves substantial win rates compared to the random and sequential baselines, with $78.33\%$, and $71.45\%$, respectively. We also outperform training with the single best RM (per RewardBench) by a $56.34\%$ win rate. We hypothesize the lower win rate of the baselines stems from the inability of these baselines to deal with conflicting signals from multiple RMs (see Fig. 5 for further analysis). The results are also applicable to the Mistral-7B model, which achieved a win rate of $58.72\%$ against the best-RM baseline and a win rate of $63.72\%$ against the sequential selection baseline. Overall, these results highlight that LASeR excels in tasks without gold labels and performs consistently well at following instructions across various user queries, showcasing its adaptability to diverse tasks.
Figure 2: Length-controlled AlpacaEval win rates comparing LASeR against baselines on the instruction-following tasks on WildChat using Llama-3-8B.
LASeR’s adaptive RM selection helps long-context understanding.
Given the cost of training long-context systems, for LongBench (Bai et al., 2023), rather than finetuning a model using RMs, we employ the selected RM to rerank generation in Best-of-$n$ sampling (see Sec. 3.3). In Table 2, we observe that LASeR consistently outperforms the baselines across tasks on Llama-3-8B and Mistral-7B except on summarization, where we achieve comparable performance. LASeR improves single-doc QA by $3.58$ F1 points over the base Llama-3-8B model and $2.64$ F1 points over random RM selection. On multi-doc QA, our approach improves performance over the Llama-3-8B and Mistral-7B models by $\approx\!4$ F1 points each, beating out random RM selection by $2.42$ F1 points on Mistral-7B. Lastly, on few-shot learning tasks, LASeR provides over $3$ points gain in F1 compared to the base model for both Llama-3-8B and Mistral-7B, surpassing the average RM performance by up to $2.4$ F1 points (on Llama-3-8B) and demonstrating its effectiveness across tasks.
@@
Table 2: LASeR outperforms baselines in long-context understanding tasks with Llama-3-8B and Mistral-7B. Sequential RM selection is not applicable in this setting as only inference is conducted. For QA and few-shot learning tasks, we report F1 scores, and for summarization we report Rouge-L.
Method Single-Doc QA Multi-Doc QA Summarization Few-shot Learning
Llama-3 Mistral Llama-3 Mistral Llama-3 Mistral Llama-3 Mistral
Base model 33.89 26.01 32.96 24.06 29.54 26.47 70.23 64.93
Best RM 35.12 28.93 35.83 27.93 34.26 30.42 71.79 68.34
Avg. RM 34.96 27.47 35.39 26.38 31.71 27.34 71.28 66.94
Random RM Selection 34.83 27.44 35.19 25.38 31.57 27.19 70.91 66.72
LASeR (Ours) 37.47 29.14 36.94 27.80 34.13 30.08 73.31 68.47

### v2

@@ 3.2 Bandit Algorithms for Adaptive RM Selection
Background: Multi-Armed Bandits.   The multi-armed bandit (MAB) problem addresses the challenge of balancing exploration and exploitation in sequential decision-making (Vermorel and Mohri, 2005; Audibert et al., 2009). The goal is to maximize cumulative MAB rewards over time by selecting arms that yield the highest MAB rewards. In order to distinguish between rewards or scores generated by RMs and the rewards used in MAB literature, we refer to the latter as “MAB rewards”. A decision-making agent in a bandit setting faces a trade-off: whether to exploit the arm with the highest known MAB reward based on past observations or explore other, less familiar arms to gather more information that might lead to even better rewards in the future. In a contextual MAB setting, the agent is also provided with additional information in the form of a context, such as the current state and input, to help inform arm selection accordingly.
LASeR uses MABs to dynamically identify the most suitable RM for each query $x_{i}$ and task through exploration while simultaneously training the LLM (since this fine-grained information is not known a priori). Pulling a previously un(der)-explored arm allows the MAB to update its information about the relevance and quality of preference pairs built using that RM via the MAB reward (discussed below). If we over-explore RMs, we might waste time on underperforming RMs, slowing down the overall training progress. On the other hand, if we over-exploit, we might prematurely focus on one RM that seems best initially but is not optimal for all queries or tasks. The exploration-exploitation nature of MAB ensures that LASeR adapts to the specific needs of each query (see Appendix B for further analysis) while improving overall performance.
RM Selection in LASeR.   LASeR uses mini-batch training for each iteration, i.e., we use MABs to select a single RM for a batch of prompts $\mathbf{x}_{m,t}$ for $t^{\text{th}}$ batch or training step of iteration $m$ (total of $T$ steps/batches in each iteration). [^3] Note that LASeR can switch between RMs at the instance level if the batch size is set to 1; however, for the sake of efficiency, we batch instances together both for LASeR and the baselines, as this reduces the computational overhead associated with loading RMs onto the GPU. We provide further discussion in Sec. A.1. Let the set of $K$ reward models (or arms) be denoted by =$\mathcal{R}=\{R_{1},R_{2},\dots,R_{K}\}$, where each $R_{k}$ corresponds to a different RM. We use the negative cumulative train loss function on the batch ($-\mathcal{L}^{m}$) at the given iteration $m$, i.e., $\mathcal{L}^{m}\!=\!\mathcal{L}^{m}_{\textsc{dpo}}+\mathcal{L}^{m}_{\textsc{nll}}$ for reasoning tasks, $\mathcal{L}^{m}\!=\!\mathcal{L}^{m}_{\textsc{dpo}}$ for other task, as mentioned in Sec. 3.1, as the MAB reward. The MAB reward is computed after the model is trained on each batch. Specifically, the MAB reward is the negative training loss (DPO), which depends on how clearly the model learns to prefer the RM’s chosen outputs over the rejected ones. A lower DPO loss or a higher MAB reward corresponds to a larger log-likelihood margin between preferred and dispreferred responses, i.e., $\log\pi_{m}(y_{i}^{w}\mid x_{i})-\log\pi_{m}(y_{i}^{l}\mid x_{i})$ for a query $x_{i}$ and preference pair $(y_{i}^{w},y_{i}^{l})$, indicating that the RM’s feedback helped the model sharpen its rankings and increase confidence in distinguishing between chosen and rejected responses. Thus, an RM that provides more informative and consistent rankings is given a higher reward compared to a less-suited RM providing uninformative rankings.444 Note that since DPO training typically begins from a supervised fine-tuning (SFT) checkpoint, the policy model already has some ability to distinguish good responses from bad ones. In Fig. 5 we show empirically that the model’s margin is typically positive, further supporting this claim.
@@
We employ LinUCB (Li et al., 2010), a contextual bandit algorithm for the arm or RM selection. We choose LinUCB because it is a contextual bandit algorithm that effectively leverages context information, making it well-suited for integration into LLM frameworks due to its ability to dynamically adapt decisions based on contextual embeddings. Additionally, LinUCB has been shown to provide a good trade-off between computational efficiency and performance (Zhou, 2015). LinUCB assumes that the MAB reward can be modeled linearly as a function of context features and computes the expected MAB reward of each arm with an upper confidence bound to ensure exploration (Garivier and Moulines, 2008, 2011). In each step $t$, we have a batch of input prompts $\mathbf{x}_{m,t}$ for which we compute sentence embeddings, using the policy model $\pi_{m}$, and use the mean sentence embedding as the context $c(t)$ to the MAB, i.e., $c(t)=\sum_{x\in\mathbf{x}_{m,t}}e_{m}(x)/|\mathbf{x}_{m,t}|$, where $e_{m}(.)\in\mathbb{R}^{d}$ yields the sentence embedding from the model $\pi_{m}$. We calculate the embedding for a prompt as the last token embedding from model $\pi_{m}$ (details in Sec. A.1). The learned parameters of the LinUCB bandit include $\hat{\theta}_{k}\in\mathbb{R}^{d}$ which represents the learned weights for the features of each reward model and $A_{k}\in\mathbb{R}^{d\times d}$ (a covariance matrix) and a bias vector $b_{k}\in\mathbb{R}^{d}$ corresponding to each arm or RM $R_{k}$. We initialize the parameters for LinUCB by randomly initializing $b_{k}$ and setting parameter $A_{k}$ to the identity matrix. Based on the LinUCB algorithm, for each batch, the selected RM $R^{\star}_{t}$ is determined by $R^{\star}_{t}=R_{j}$ such that:
$j=\arg\max_{k\in[1,K]}\left(c(t)^{\top}\hat{\theta}_{k}+\alpha\sqrt{c(t)^{\top}A_{k}^{-1}c(t)}\right),$ (1)
@@
We emphasize that the effectiveness of our approach is not tied to particular bandit algorithm but rather from the exploration-exploitation trade-off of MAB framework. LASeR is agnostic to the bandit algorithms and can incorporate alternatives to LinUCB. Indeed, we provide a comparison of LinUCB and other bandit algorithms in Appendix C.
@@ 3.3 LLM and Bandit Training in LASeR
LASeR with Best-of-$n$ Sampling.   For settings where finetuning the LLM is not desirable or feasible, LASeR can also be applied to learn the MAB parameters without training the LLM. Rather than fine-tuning the model with preference data, we employ best-of-$n$ sampling (Lightman et al., 2023; Sun et al., 2024), where multiple responses are generated, and the best one is selected based on the RM. The bandit parameters are then updated using equation (2), with the MAB reward calculated as the negative normalized NLL loss on the train data. This updated bandit can subsequently be used for inference on the test set. This approach is particularly useful for long-context understanding tasks, where training would be too computationally intensive (see setting in Sec. 4.2).
@@ 4.1 Experimental Setup
Models.   We conduct our experiments on the Llama-3-8B base (AI@Meta, 2024) model. We present additional results with Mistral-7b-v3-Instruct (Jiang et al., 2023) and Qwen2.5-32B (Qwen Team, 2024a) in Appendix B. For training, all models are fine-tuned using Low-Rank Adaptation (LoRA) (Hu et al., 2021) for efficiency. For both training and inference, we do 0-shot prompting and sample $n\!=\!30$ responses per prompt with temperature $0.8$ (see Sec. A.1 for more details).
@@
Reward Models.   We select $K=4$ strong 7B RMs from RewardBench (Lambert et al., 2024), which include Zephyr-7B-Alpha, Qwen1.5-7B-Chat, Eurus-7B-KTO, and OLMo-7B-Instruct. Following the pipeline outlined in Lambert et al. (2024), for these models, we compute the reward for each response as the log-likelihood of the RM for that response (see Secs. A.1 and D for more details and discussion about the RM choices).
Datasets and Metrics.   Our experiments cover a range of tasks and datasets (see Sec. A.1):
@@
  • Reasoning: Evaluating reasoning abilities is crucial for testing the model’s capacity to handle complex, multi-step tasks and has presented a challenge to iterative preference optimization methods (Yuan et al., 2024b; Chen et al., 2024b). We train and evaluate on StrategyQA (Geva et al., 2021), MMLU (Hendrycks et al., 2021b, a), and GSM8K (Cobbe et al., 2021).
@@
  • Instruction-Following: We further evaluate our method on heterogeneous tasks without gold labels. We use user prompts from WildChat dataset (Zhao et al., 2024), which contains a collection of natural user-chatbot interactions. This dataset has five primary categories of instruction-following prompts: creative writing, analysis, coding, factual information, and math reasoning. Due to computational constraints, we randomly subsample 5K prompts from each category for model training. We compare models trained with LASeR against baselines (described below) using length-controlled AlpacaEval (Dubois et al., 2024) that pairs responses from two different LLMs and uses GPT-4o as a judge to pick the winner, accounting for the length of both responses.
@@
  • Long-Context Understanding: As finetuning LLMs on long-context inputs is computationally intensive, we demonstrate the effectiveness of LASeR using Best-of-$n$ sampling on LongBench (Bai et al., 2023), which consists of multiple tasks, such as single-document QA, multi-document QA, summarization, and few-shot learning. For summarization, we report Rouge-L (Lin, 2004), whereas for the rest, we report the F1 score.
Baselines.   We compare LASeR against two baseline categories reflecting different kinds of methods for using RMs in LLM training: single RM selection and multi-RM ensemble baselines. For single RM selection baselines, we evaluate against the following baselines:
@@
  • Best RM: From our collection of RMs, we pick the RM that corresponds to the best overall score on RewardBench (Lambert et al., 2024): Zephyr-7B-Alpha. We use this single RM during training (c.f. Sec. 3.1). This baseline reflects the performance gain a user could expect when selecting the best RM from a leaderboard without knowing a priori how it generalizes to a particular task.
@@
  • Avg. RM: Here, we perform single RM training over all the RMs in the pool and report the average performance. A comparison with this baseline represents an expected gain from a randomly picked RM from a leaderboard; however, note this baseline consists of multiple models averaged together.
@@
  • Random RM Selection: In this baseline, we randomly sample a single RM from the set of RMs (from a uniform distribution) for each training batch in every iteration.
@@
  • Sequential RM Selection: In training, this method explores RMs sequentially and based on a set order in each iteration to examine their impact on model training, demonstrating that, instead of optimizing with all RMs, LASeR can adaptively select the best RM for each batch of samples.
  • Classifier Selection: To compare against a context-sensitive baseline that does not use an MAB, we train a $K$-way classifier to perform RM selection using data from RewardBench (see Sec. A.1). Specifically, for each query and RM, we compute the RM’s score of the annotated preferred and dispreferred response. The RM that assigns the correct preference ordering with the highest difference between the scores of the preferred and dispreferred responses is chosen as the RM label and used to train the classifier. While training the LLM, for each input $x_{i}\in\mathcal{D}$, we select the RM for building preference pairs based on this trained classifier.
Additionally, we compare against the following RM ensemble baselines:
@@
  • RM Score Ensemble: We generate multiple responses for each query, which are then scored using each RM, and the preference dataset is created by averaging the scores across all RMs (following Coste et al. (2023)); thus, comparing LASeR with using all RMs simultaneously.
  • RM Agreement Ensemble: Because ensembling scores through averaging is sensitive to the absolute scores produced (which may differ between RMs), we follow Wang et al. (2024d) in ensembling through ranking and agreement. Specifically, we generate 32 responses for each query and sample 100 pairs for each. We score each pair with RMs, constructing a preference dataset by choosing the 10 pairs with the highest agreement of preference rankings across RMs.
  • Online RM Ensemble: Since uniformly-ensembled RM scores might fail to prevent the LLM from exploiting an underperforming RM, we further compare our method to an additional ensemble baseline that uses an online learning approach. Instead of setting uniform weights for each RM as in RM Score Ensemble, each RM is assigned a learned weight; these weights are dynamically updated after each training batch using the multiplicative weights algorithm (Arora et al., 2012).
Conceptually, among the single RM selection methods, the best RM baseline serves as an “exploit-only” setting that exploits the best available RM based on aggregate RewardBench scores. On the other hand, the random and sequential selection baselines are “explore-only” in that they pick a new RM either randomly or via a predefined sequence, irrespective of the performance of each arm (RM). LASeR and the classifier approach represent two methods for selecting a single optimal RM for each instance; however, we note that the classifier approach depends on a fine-grained, in-distribution dataset of queries paired with their corresponding suitable RMs for annotating preferences, which is often absent in practice. In our experiments, we demonstrate the benefits of balancing the exploration-exploitation trade-off compared to “exploit-only” and “explore-only” approaches. Additionally, different RM ensemble baselines employ various methods for combining RMs while utilizing them simultaneously. Furthermore, we analyze potential conflicting signals among RMs to highlight the advantages of our framework over RM ensemble baselines (Sec. 5). The settings for training LASeR and baselines are provided in Sec. A.1.
@@ 4.2 Main Results
LASeR achieves the best average accuracy on reasoning tasks.   Table 1 shows our method consistently outperforms the baselines across multiple benchmarks, particularly in the StrategyQA and GSM8K tasks. For example, LASeR improves approximately $2\%$ absolute accuracy over the sequential baseline on both GSM8K and StrategyQA. In cases where the best RM is not known beforehand, LASeR surpasses the performance of the average RM by $2.7\%$ and the RM Score Ensemble for each instance by $2.67\%$ (in accuracy averaged over the three tasks). Moreover, this lower performance by the RM Score baseline is not purely due to variance in the scores: LASeR also surpasses the RM Agreement Ensemble and RM Online Ensemble by $0.91\%$ and $2.27\%$, respectively. Overall, LASeR provides consistent results while the underlined second-place models show inconsistent performance across datasets. These results also emphasize the benefit of LASeR, as it eliminates the need to choose a different RM in advance or ensemble multiple RMs.
LASeR beats baselines at instruction-following.   Often, LLMs are used by large numbers of people with a diverse set of queries, goals, and intentions, and their preferences vary based on the underlying query. To demonstrate the effectiveness of LASeR in such settings,
Table 1: Performance on reasoning benchmarks on Llama-3-8B. The baselines also include supervised fine-tuning on human-written responses (SFT) as a reference for performance without preference optimization. The highest accuracy is shown in bold, and the second-highest accuracy is underlined. LASeR yields the highest average accuracy.
Method StrategyQA GSM8K MMLU Avg.
SFT 80.41 69.43 65.66 71.83
Best RM 84.29 73.16 67.15 74.87
Avg. RM 82.62 71.57 66.67 73.62
Random RM Selection 84.37 71.99 67.85 74.74
Seq. RM Selection 83.90 72.94 68.02 74.95
Classifier Selection 83.13 72.73 67.96 74.60
RM Score Ensemble 82.96 70.94 67.04 73.65
RM Agree. Ensemble 84.03 73.85 68.35 75.41
RM Online Ensemble 83.25 72.04 66.85 74.05
LASeR (Ours) 85.96 74.75 68.24 76.32
@@
we compare the instruction-following performance in Fig. 2, i.e., AlpacaEval win rates, of LLMs trained using LASeR with the baselines using WildChat. Specifically, LASeR achieves substantial win rates compared to single RM selection baselines such as random and sequential selection, with $78.33\%$, and $71.45\%$, respectively. We also outperform training with the single best RM (per RewardBench) by a $56.34\%$ win rate. Compared to RM ensemble baselines, LASeR achieves $72.69\%$ and $67.88\%$ win rates over RM Score Ensemble and RM Online Ensemble, respectively. We hypothesize the lower win rate of the baselines stems from the inability of these baselines to deal with conflicting signals from multiple RMs (see Fig. 3 for further analysis). Overall, these results highlight that LASeR excels without gold labels and performs consistently well at following instructions across various user queries, showcasing its adaptability to diverse tasks.
LASeR’s adaptive RM selection helps long-context understanding.   Given the cost of training long-context systems, for LongBench (Bai et al., 2023), rather than finetuning a model using RMs, we employ the selected RM to rerank generation in Best-of-$n$ sampling (see Sec. 3.3).
@@
Table 2: LASeR outperforms baselines in long-context understanding tasks with Llama-3-8B. Sequential RM selection is not applicable in this setting, as only inference is conducted. For QA and few-shot learning tasks, we report F1 scores, and for summarization, we report Rouge-L.
Method SiQA MuQA Sum FS
Base model 33.89 32.96 29.54 70.23
Best RM 35.12 35.83 34.26 71.79
Random RM Selection 34.83 35.19 31.57 70.91
Classifier RM Selection 34.42 34.24 32.41 70.58
RM Score Ensemble 34.51 35.52 32.38 70.34
RM Agree. Emsemble 35.79 35.41 33.19 72.15
RM Online Ensemble 34.69 35.80 32.89 70.51
LASeR (Ours) 37.47 36.94 34.13 73.31
In Table 2, we observe that LASeR consistently outperforms the baselines across tasks. LASeR improves single-doc QA (SiQA) by $3.58$ F1 points over the base Llama-3-8B model and $2.64$ F1 points over random RM selection. On multi-doc QA (MuQA), our approach improves performance over the Llama-3-8B base model by $\approx\!4$ F1 points. Furthermore, on few-shot (FS) learning tasks, LASeR provides over $3$ points gain in F1 compared to the base model, surpassing the average RM performance by up to $2.4$ F1 points and demonstrating its effectiveness across tasks. Lastly, Table 2 demonstrates that LASeR consistently outperforms the RM Score Ensemble baseline across different long-context tasks (except for the Summarization task where LASeR is comparable to Best RM), e.g., a $\approx\!3$ F1 point boost on single-doc QA and few-shot learning tasks.
Figure 2: Length-controlled AlpacaEval win rates comparing LASeR against baselines on WildChat instruction-following tasks using Llama-3-8B. The top row shows comparisons against single RM selection methods, while the bottom row shows comparisons against multi-RM ensemble methods.